OpenStudy (anonymous):

Given the equation 2x + 4 = 4x - 2, select the reasoning that correctly solves for x. Add 2, subtract 2x, then divide by 2. Add 2, subtract 4x, then divide by -2. Subtract 4, subtract 2x, then divide by -2. Subtract 4, subtract 4x, then divide by 2.

OpenStudy (anonymous):

add 2, subtract 2x, then divide by 2 2x + 4 = 4x - 2 (add 2 to both sides) 2x + 4 + 2 = 4x 2x + 6 = 4x (subtract 2x from both sides) 6 = 4x - 2x 6 = 2x (divide by 2) 3 = x